10.A car accelerates uniformly from rest to a speed of 30.0 mi/h in 11.0 s.

(a) Find the distance the car travels during this time.

 m


(b) Find the constant acceleration of the car.

 m/s2

1
Expert's answer
2022-06-03T15:25:23-0400

Given:

"v=30\:\rm mi/h=13.4\;\rm m/s"

"t=11.0\:\rm s"

(a)

"d=\frac{v+v_0}{2}t=\frac{0+13.4}{2}*11.0=74\;\rm m"

(b)

"a=\frac{\Delta v}{\Delta t}=\frac{13.4}{11.0}=1.22\:\rm m/s^2"